# Low energy phonons in single crystal ZrW2O8

### v1

@@ III.2 DFT calculations
The DFT calculations were performed at a lattice parameter of 9.215Å, and we applied a global re-scaling of the calculated phonon energies of 95% to give best agreement with the INS measurements  [^2] This was done by doing a least squares fit of the neutron spectra with the DFT-calculated spectra, varying only the scale factor to give the best agreement. Coincidentally, the relative difference in lattice parameter between that used for the DFT and that observed experimentally at low temperature is approximately the same as the experimentally observed negative thermal expansion between 100 K and 400 K [9]. We show in fig. 8 the effect of the lattice parameter on the computed phonon spectra below 6 meV (the energy range covered in our INS experiments). As would be expected from the Grüneisen parameters reported in the literature, the low energy phonons are mostly found to harden with increasing lattice parameter. However, the scaling is not uniform, as can be seen in the inset which shows the trajectory along Z – M2 – Y. Here it is clear that the first optic mode (labelled 2 and 4 for the simulation with $a=9.259$  (the optimized lattice parameter in the DFT calculations) and $a=9.215$  respectively) is significantly more dispersive between Z and M2 for the larger lattice parameter. This also illustrates why we chose to compare the INS data with the calculations for the unit cell with $a=9.215$  rather than $a=9.259$. A global re-scaling of the latter results in a worse agreement with the measurement than the former.
@@
Figure 8: Comparison of dispersion from DFT calculations with different lattice parameter. The reciprocal space path is the same a described in fig. 2. The solid red curve is for calculations with lattice parameter $a=9.259$(zero stress, optimized lattice parameter), and the dashed blue curve is for calculations with $a=9.215$. The inset shows a zoomed in section of the dispersion between 3 and 6 meV. The modes labelled 1, 2 and 5 correspond to calculations with $a=9.215$, and the modes labelled 3, 4 and 6 correspond to calculations with $a=9.259$. The symbols at high symmetry wavevectors correspond to the fitted peak positions for the lower temperature datasets shown in fig. 7, with diamonds for the G-point, stars for the Z-point, pentagons for the M2-point and a circle at the R-point.
One might naturally ask if further compression of the lattice in the calculations could yield yet better agreement. In fig. 9 we show calculations with $a=9.215$  and $9.210$. Here we see that certain modes, for example on the path Y – R – G, have softened considerably, such that an overall re-scale of energies would once again not yield good agreement with the INS data. What this seems to illustrate is that the calculations which agree best with the INS data are those which are on the edge of dynamical stability. Indeed, as noted in sec. II.1, when the lattice was compressed further in the DFT calculations, a significant number of soft modes appeared, consistent with this interpretation. With further compression negative energy modes appear, indicating an eventual change in the ground state structure.
@@
Figure 9: Comparison of dispersion from DFT calculations with lattice compressed further. The reciprocal space path is the same a described in fig. 2. The solid cyan curve is for calculations with lattice parameter $a=9.210$, and the dashed blue curve is for calculations with $a=9.215$.
@@
Fig. 10 shows the partial phonon density of states (PDOS) computed from DFT, broken down into the contributions from each of the constituent atomic species. The advantage of computing this property is that it is intrinsically averaged across all of reciprocal space and hence offers some insight into the combined effect of many different phonon modes to the material’s thermal properties. The disadvantage, as already noted, is that the averaging means that information about particular phonon modes is lost.
In the energy range probed by this experiment, the PDOS has approximately equal weight from tungsten and oxygen, but very little contribution from zirconium. This indicates that the zirconium ions are much more fixed in position, somewhat characteristic of rigid-unit-modes. Further decomposition of the PDOS down to the individual atom level (see supplementary information [32]) shows that the contribution from individual tungsten ions is larger than for individual oxygen ions by about a factor two to three at low energies, but the oxygen ions are more numerous so sum up to give a comparable contribution to the PDOS.
@@
Figure 12: Slices through the $\mathbf{Q}$ energy plane, following the same scheme as fig. 3. The colormaps show the neutron scattering intensity calculated using DFT, decomposed into the contributions from each individual element. Panel (a) shows the contribution from zirconium, panel (b) the contribution from tungsten, and panel (c) the contribution from oxygen. Note that the intensity scale is the same as that in fig. 3 and is the same for all three panels.

### v2

@@ III.2 DFT calculations
The DFT calculations plotted in figures. 24 were performed at a lattice parameter of 9.215Å, and we then reduced all calculated phonon energies by 5% to give best agreement with the INS measurements  [^2] This was done by doing a least squares fit of the neutron spectra with the DFT-calculated spectra, varying only the scale factor to give the best agreement.
@@
We show in fig. 8 the effect of the lattice parameter on the computed phonon spectra below 6 meV (the energy range covered in our INS experiments). As would be expected from the Grüneisen parameters reported in the literature, the low energy phonons are mostly found to harden with increasing lattice parameter. However, the scaling is not uniform, as can be seen in the inset which shows the trajectory along Z – M2 – Y. For instance, the two lowest energy modes for $a=9.210$  and $a=9.215$  are relatively flat in energy, whereas the equivalent modes for $a=9.259$  are much more dispersive. Furthermore, one can see that at the M2 point the highest energy mode decreases by $\sim 0.2$ meV on compression of the lattice from $a=9.259$  to $a=9.215$  ($\Delta a=-0.044$) and decreases by even more, $\sim 0.3$ meV, for further compression $\Delta a=-0.005$. As noted above, we compared the INS data with the calculations for the unit cell with $a=9.215$  rather than $a=9.259$, which is the zero-stress optimized unit cell dimension in the DFT, or the cell with $a=9.210$. Globally rescaling the phonon energies in either of the latter two sets of calculations resulted in worse agreement with the neutron data.
@@
Figure 8: Comparison of dispersion from DFT calculations with different lattice parameters. The reciprocal space path is the same a described in fig. 2. The solid red curve is for calculations with lattice parameter $a=9.259$  (zero stress, optimized lattice parameter), the dotted blue curve is for calculations with $a=9.215$, and the dash-dotted cyan curve is for calculations with $a=9.215$. The inset shows a zoomed in section of the dispersion between 3 and 6 meV. The symbols at high symmetry wavevectors correspond to the fitted peak positions for the lower temperature datasets shown in fig. 7, with diamonds for the Gamma - point, stars for the Z - point, pentagons for the M2 - point and a circle at the R - point.
As noted, a small compression of the lattice from $a=9.215$  to $a=9.210$  yields a rather large change in mode energies, suggesting that around these lattice parameters the calculations are on the edge of dynamical stability. Indeed, when the lattice was compressed further in the DFT calculations (not shown), a significant number of soft modes appeared, consistent with this interpretation. With further compression negative energy modes appear, indicating a change in the ground state structure. Given this non-linear behavior, to make a quantitative prediction for the effect on the phonon energies of temperature-induced lattice compression in this quasi-harmonic approximation, it makes most sense to compare the calculations with $a=9.259$  and $a=9.215$, rather than those with $a=9.215$  and $a=9.210$. The relative difference in the former two lattice parameters is 0.48%, which is quite close to the experimentally observed change in lattice parameter on warming from 100 K to 400 K (0.37%). Thus, one can take the changes in the phonon dispersion between these two calculations as representative of the changes in the phonon dispersion over the experimentally probed temperature range, based on the Grüneisen model alone. The results of this procedure at various high symmetry wavevectors are presented in table 1.
Table 1: Comparison of the position of the lowest energy peak in the measured data at 100 K and 400 K, and the expected change in peak position based on DFT alone (i.e. quasi-harmonic approximation), and the computed change in peak position from the work of Gupta et al [30], decomposed into the contributions from implicit and explicit anharmonicity, from a phenomenological model.
Wavevector Lowest energy peak at 100 K (meV) Lowest energy peak at 400 K (meV) DFT shift (meV) [this work] Calculated shift [implicit + explicit] (meV) [30]
G 5.10(3) 5.12(5) -0.55 -0.07 [-0.22 + 0.15]
Z 3.50(5) 3.70(10) -0.40 0.08 [-0.14 + 0.22]
M 3.77(4) 3.58(7) -0.78 0.06 [-0.36 + 0.42]
R 4.45(2) 4.43(5) -0.80 -0.77 [-0.39 - 0.38]
The DFT calculations presented so far correspond to the lattice contracting or expanding under pressure / stress, rather than warming / cooling. In reality, the warming will result in an increased amplitude of atomic vibrations, which is not accounted for in the quasi-harmonic approximation. Anharmonic effects may become more significant at higher temperatures. This point has been discussed extensively by Mittal et al [30, 12], who have calculated these effects for a few individual modes at particular wavevectors in ZrW2O8 . Some attempt has been made to quantify the effect more broadly using Raman spectroscopy data [17, 18], which has indicated that as well as having large negative Grüneisen parameters, the lowest energy modes also have the largest anharmonicity.
The effects of temperature on phonons can be decomposed into two parts. On the one hand a change in lattice parameter with temperature often gives rise to a changed bond length between atoms hence a change in phonon energy. This is sometimes referred to as implicit anharmonicity. In conventional materials, increased temperature gives longer and weaker bonds, and hence softened phonons. On the other hand, as temperature increases phonon-phonon interactions become increasingly important, sometimes also referred to as explicit anharmonicity, and renormalize the phonon energies. This often yields phonon modes that harden slightly with increasing temperature [16]. However, in practice in conventional materials this effect is often much smaller than the softening due to thermal expansion and so is masked.
In table 1 we compare the positions of peaks in the INS signal at high symmetry wavevectors as a function of temperature, with the expected shift due to a change in lattice parameter alone from the DFT. Additionally, we compare these results to phenomenological calculations by Gupta et al [30] that decompose the shift into components arising from implicit and explicit anharmonicity. With the exception of the calculation at the R–point, the changes in peak position from the latter are much smaller than those predicted from the DFT alone, and are closer in magnitude to the effects seen in the INS measurements.
III.3 Machine learned potentials and molecular dynamics calculations
[…]
In order to compare the MD results to the INS data, dynamical structure factor calculations were performed with the lattice parameters fixed to the time-averaged values at each temperature. Gaussians were then fitted to the peaks in energy scans up to 6 meV at several high symmetry wavevectors to quantify the temperature-dependent shifts in these modes. The results are shown in fig. 9, with the energy scans shown in the appendices figs. 1417. The majority of the modes harden slightly on warming, with the exception of the highest energy one at the M point ($\mathbf{q}=(0.5,0.5,0)$) which softens slightly. We note that between 100 K and 400 K the calculated magnitude of mode hardening of $\sim 0.1$ meV is similar in magnitude to what is seen in our INS measurements. But between 0 K and 100 K a much greater mode hardening is predicted, notably with the modes at $\mathbf{q}=(0,0.5,0)$ and $\mathbf{q}=(0.25,0.5,0)$ hardening on warming by $>0.5$ meV, which is not observed in the measurements. This may, for example, be due to the precise choice of functional used for the DFT training data on which the interatomic potentials were based. We note that the DFT calculations presented in sec. III.2 had to be fine-tuned for best agreement with the data, whereas the DFT training data for the MD calculations were not adjusted in this way.
Figure 9: Temperature dependence of phonon modes below 6 meV computed using MD with machine-learned potentials. Red circles indicate the modes at $\mathbf{q}=(0.5,0.5,0)$ (M1 in the notation of fig. 2), green squares the modes at $\mathbf{q}=(0.25,0.5,0)$, and blue circles the modes at $\mathbf{q}=(0,0.5,0)$ (Y in the notation of fig. 2).
IV Discussion and Conclusions
From the results and analysis presented, we can see that there is very good agreement between the phonon dispersion relations and structure factors measured in our inelastic neutron scattering experiments, and those computed from DFT in the zero-temperature quasi-harmonic approximation, albeit with a compressed lattice compared to the optimized value. The agreement is good throughout reciprocal space, which is important in the context of the view that there is no single phonon mode responsible for NTE in ZrW2O8, but many. This therefore supports the usefulness of the DFT results in interpreting the physics of NTE in ZrW2O8. We do not observe any significant change in the neutron structure factors as a function of temperature, aside from some modes broadening on warming, so we can be confident that the phonon eigenvectors do not change significantly with temperature. The fundamental nature of the atomic motions thus do not change with temperature, which is consistent with DFT.
On the other hand, the DFT predicts softening of most phonon modes on compression of the lattice, and with NTE such a compression happens on warming. But we observe that the majority of phonons do not soften, indeed some harden by an amount ($\sim 0.1$ meV) that is just discernible with our instrumental resolution. This is because DFT’s quasi-harmonic calculations only capture the lattice compression on warming, but do not capture the full anharmonicity, which plays an increasing role with increasing temperature. Finite temperature effects can be included in MD simulations, which in our case used potentials generated through machine learning applied to DFT. Our MD simulations showed that the low energy modes at a few high-symmetry wavevectors do indeed harden slightly on warming. A larger hardening than what is observed was predicted between 0 K and 100 K, but as already noted this may be due to the specific choice of functional for the DFT training data for the generation of the interatomic potentials.
Given that the DFT calculations give good phonon eigenvectors when compared to experiment, and the MD calculations show the right behavior of the phonon eigenvalues as a function of temperature when compared to experiment, it is instructive to examine the underlying atomic motions that are common to both to help us understand NTE in ZrW2O8 . Such analysis has been done separately for the DFT [28] and MD [38] respectively. The former analysis compares the DFT to a set of simplified flexibility models, and concludes that NTE arises due to a tension effect involving nearly rigid Zr–O and W–O bonds, with low energy phonons likely involving quasi-rigid small whole-body rotations of the WO4 tetrahedra. The analysis of the MD shows that with increasing temperature the Zr–O and W–O bond lengths indeed change very little, with a small positive expansion, but the W–O–Zr bond angle changes significantly, that could also be described as a tension effect. Our INS data are consistent with both sets of analysis, and both point towards the same basic mechanism of NTE.
Acknowledgements.
We are grateful to D. Voneshen (ISIS) and R. Fair (STFC-SCD) for their advice and assistance on the use of Euphonic. We are indebted to M. Dove (QMUL) for his insightful and informative comments. Raw data are available via ref. 39 and processed data via ref. 40. We thank the Institut Laue-Langevin for the allocation of beamtime for the experiments reported here. Computing resources were provided by the STFC Scientific Computing Department’s SCARF cluster. AID gratefully acknowledges funding from the Ada Lovelace Centre (ALC) at the Science and Technology Facilities Council (STFC) and from the Norwegian Research Council under grant ”Neutron Scattering and Atomistic Simulations for a SUPERior Understanding of SUPERionic Conduction” (grant no. 344062).
Appendix I: Methods
A: Inelastic neutron scattering
ZrW2O8 crystallizes in the $P2_{1}3$ cubic space group. The single-crystalline samples for these experiments were grown as described previously [41]. The sample used was composed of one crystal of mass 2.4 g with a mosaic of $\sim 0.3$ degrees, which is substantially less than the divergence of the incident neutron beam in the inelastic neutron scattering experiments described below, and hence unimportant for the experimental resolution. The lattice parameter was measured to be 9.176 Å  at 2 K, 9.159 Å  at 100 K, 9.1389 Å  at 300 K, and 9.125 Å  at 400 K during the alignment procedure for the inelastic scattering measurements described below. These values are consistent with those reported in the literature for powder samples [9].
The time of flight (ToF) neutron scattering experiments were performed on the IN5 spectrometer at the Institut Laue-Langevin [32]. Data were collected with incident neutron energy $E_{i}=8.24$ meV, with the final chopper spun at 16,000 r.p.m. to give a resolution at the elastic line of 0.27 meV (full-width half-maximum, FWHM). The sample was aligned so that the $(H,H,0)$ and $(0,0,L)$ reciprocal space directions were in the horizontal plane of the instrument. The data were collected by scanning the sample orientation about the vertical axis in 0.5 degree steps over a range of 161 degrees, starting from the $(1,1,0)$-axis perpendicular to the incident neutron beam. Data were collected at temperatures of 100 K and 400 K, with these temperatures chosen in order to increase the signal via the Bose-Einstein population factor while maintaining a large temperature difference between measurements, but avoiding the structural phase transition at $\sim 425$ K [9]. Additional data were collected with the sample aligned with the $(H,0,0)$ and $(0,K,0)$ directions in the horizontal plane of the spectrometer, with the sample orientation scanned over a range of 88 degrees in 1 degree steps. These data were collected with $E_{i}=10.43$ meV and a final chopper speed of 16,000 r.p.m. giving a FWHM energy resolution at the elastic line of 0.35 meV. Data with this configuration were collected at temperatures of 2 K and 300 K.
The raw data were processed (units conversion, normalization of intensities, etc.) using the Mantid software package [42]. Further processing of the data, and all subsequent visualization and analysis, including convolution of calculated spectra with the instrumental resolution, were performed using the Horace software package [43]. Calculation of the inelastic neutron scattering cross-section ($S(\mathbf{Q},\omega)$) from the DFT-calculated phonons was done using the Euphonic software package’s Horace plugin [44, 45].
B: Density functional theory
The phonon dispersion was calculated using Density Functional Perturbation Theory (DFPT) [51], with a $3\times 3\times 3$ Monkhorst-Pack grid of phonon wavevectors and Fourier interpolation of the resulting dynamical matrices, in order to sample the entire Brillouin zone. Additional calculations were performed for fixed unit cells of smaller dimension (9.23, 9.22, 9.215, 9.21 and 9.16 Å), at each of which the structure was relaxed and phonons calculated in the same way as for the fully relaxed structure. The calculated stress for each of these cell dimensions was 0.81, 1.13, 1.29, 1.45 and 3.20 GPa respectively. In the first four of these cases no negative energy phonon modes were observed, indicating that the computed structure remained dynamically stable. For the calculation with a lattice parameter of 9.16 Å a large number of negative energy modes were observed, indicating that for such a compressed lattice the structure becomes unstable. From this point onward when we refer to these DFT calculations we mean those with $9.21\leq a\leq 9.23$.
When comparing the computed dispersion with that measured experimentally, an empirical scale factor was applied to the eigenvalues (energies) but not the eigenvectors (structure factors). This was determined by finding the simulation that agrees most closely with the data after such a uniform re-scale: the structure factors from the simulations with different lattice parameters were fitted to the neutron data with a single adjustable parameter (the energy scale). For ZrW2O8 the best overall agreement was found for the case of $a=9.215$   and a scale factor $\sim 0.95$. This approach was used because changing the lattice parameter for the simulation does not affect the eigenvalues uniformly.
C: Machine learned potentials and molecular dynamics calculations
Calculations were performed using the automated potential development (APD) workflow [34], which is open-source and available from gitlab [53]. APD automates all DFT calculations, potential optimizations and potential simulations, as well as any subsequent cycles of active-learning that may be required to refine the potential. In addition APD also generates various properties at the DFT and potential level. Note that the DFT calculations performed by APD are distinct from those used elsewhere.
CASTEP with optimized norm-conserving pseudo-potentials [47, 54] was used as the DFT engine for APD. Calculations were performed using two generalized gradient approximation exchange-correlation functionals, both PBE [55] and the PBEsol functional [56]. For all except the phonon calculations a Monkhorst–Pack grid of 2$\times$2$\times$2 (with 44 atom cell) was used. For configuration generation, a relatively low cutoff of 520 eV was used. Molecular dynamics (MD) simulations were run at 333 K, 666 K and 1000 K, each for 2000 configurations with a timestep of 7 fs. Uncorrelated snapshots (separated by $\sim$ 15 configurations) were then recalculated using a cutoff of 1000 eV, ensuring an accuracy of ±$\pm 1$ meV/atom in energies. This cutoff, and the k-points grid, were automatically optimized by APD to achieve this accuracy.
Moment tensor potentials (MTPs) [57] were then optimized using the MLIP2 software package [58], and fitted to energies, forces and stresses (EFS) from the DFT simulations. Five distinct potentials (MTP1-5) were independently optimized. Once optimized, MTP1-5 were used to drive LAMMPS (Large-scale Atomic/Molecular Massively Parallel Simulator) molecular dynamics simulations [35]. LAMMPS jobs were performed up to 1000 K in 100 K increments, each for 10 ps using the Nosé-Hoover thermostat and barostat. From these simulations, configurations for which EFS predictions are unreliable were identified using MLIP’s MaxVol algorithm, using an extrapolation grade of 3 as a threshold. These configurations were then recalculated using DFT, and subsequently used to reoptimize potentials. The procedure was repeated until no configurations exceed an extrapolation grade of 3 during the LAMMPS simulations conducted at all temperatures. This is handled separately (and automatically) for each potential, so that each undergoes its own active-learning cycle.
As an additional post-processing step, dynamic structure factors were computed for a selection of q-points and across a range of temperatures. Dynamical structure factors were calculated using the Dynasor open-source framework [59]. Calculations were performed at high-symmetry points $Y=(0,0.5,0)$, $M=(0.5,0.5,0)$, and the midpoint between these points. A total of 200 runs, each lasting 1 ns, were performed for each temperature 100 K, 200 K, 300 K and 400 K. Each run included a 200 ps equilibration with the Number Volume Temperature (NVT) ensemble, followed by a 60 ps equilibration and a 400 ps production run with the Number Temperature Energy (NTE) ensemble. Structure factors were calculated from each run and subsequently averaged.
Appendix II: Distinguishing phonon modes via their structure factors
The coherent inelastic neutron scattering cross-section from single-phonon excitations (creation) is given by:
$\left(\frac{d^{2}\sigma}{d\Omega dE_{f}}\right)$ $=$ $\frac{k_{i}}{k_{f}}\frac{\sigma_{\rm{coh}}}{4\pi}\frac{(2\pi)^{3}}{v_{0}}\frac{1}{2M}e^{-2W}\times$ (1)
$\sum_{s}\sum_{\mathbf{G}}\frac{(\mathbf{Q}\cdot\mathbf{e}_{s})^{2}}{\omega_{s}}\langle n_{s}+1\rangle\times$
$\delta(\omega-\omega_{s})\delta(\mathbf{Q}-\mathbf{q}-\mathbf{G})$
where $k_{i}$ and $k_{f}$ are the incident and final neutron wavevectors, $\sigma_{\rm{coh}}$ is the coherent neutron cross-section of the material being studied, $v_{0}$ is the sample volume, $M$ the mass of the sample, $W$ the Debye-Waller factor, $\mathbf{Q}=\mathbf{G}+\mathbf{q}$ the excitation wavevector, composed of the Brillouin zone vector $\mathbf{G}$ and the reduced phonon wavevector $\mathbf{q}$, $\mathbf{e}_{s}$ is the phonon eigenvector for mode $s$ with frequency $\omega_{s}$, and $\langle n_{s}+1\rangle$ is the Bose-Einstein population factor. The key term is the one involving the dot product of the wavevector $\mathbf{Q}$ and the phonon eigenvector $\mathbf{e}_{s}$, which implies that the scattering cross-section is largest when $\mathbf{Q}$ and $\mathbf{e}_{s}$ are parallel, and zero when they are perpendicular. By measuring with many different orientations of $\mathbf{Q}$ all of the phonons will give rise to a signal in the INS measurement. Such an approach permits modes which may be similar in energy but with different eigenvectors to be resolved.
We note that single crystal inelastic neutron scattering measurements of ZrW2O8 have been reported [60], but with data collected at only a few high symmetry points in a small number of Brillouin zones, and without sufficient energy resolution to discern subtle shifts with a high degree of certainty. Inelastic neutron scattering experiments with single crystal samples have been performed on the structurally simple NTE materials ReO3 [61] and Cu2O [62] in the 1970s and 1980s, again covering only a few high symmetry points in a small number of Brillouin zones. More recently, as time-of-flight neutron spectrometers have developed, the multi-Brillouin zone measurement technique used in this study has been employed on structurally simple compounds exhibiting anomalous thermal expansion, e.g. Cu2O [63] and NaBr [64], permitting a detailed analysis of the phonon spectra.
Appendix III: Calculated phonon dispersion and density of states
In figure 10 we compare calculations for the PBE and PBEsol functionals with the same lattice parameter of 9.215 Å. It can be seen that using the results from the PBEsol functional result in slightly higher phonon energies for the same lattice parameter. As with the comparison between different lattice parameters using the PBE functional, shown in the main text, we note that at certain trajectories and points there are more obvious differences. For example, the inset to fig. 10 shows that the modes at the R point are significantly closer together in energy for the PBEsol functional than they are for PBE.
@@ Acknowledgements.
Figure 10: Comparison of dispersion from DFT calculations with different choices of functional, with the same lattice parameter of $a=9.215$. The reciprocal space path is the same as described in fig. 2 of the main text. The solid black curve is for calculations with the PBEsol functional and the dashed blue curve is for calculations with the PBE functional (the same as the dashed blue curve in fig. 7 of the main text). The inset shows a zoomed in section of the dispersion between 3 and 6 meV.
We therefore conclude that for analysing the INS data, calculations with either the PBE or the PBEsol functional are valid. On this basis we chose not to proceed with further simulations using the PBEsol functional (e.g. full geometry and lattice parameter optimization) since there would likely be little to be gained over the PBE calculations already performed.
Qualitatively, it would seem that the modes which are most sensitive to perturbations to the choice of psuedopotential are also the modes that seem to be most difficult for any of the DFT calculations to reproduce, i.e. those modes between about 3 and 5 meV. These are also the modes with the largest magnitude mode Grüneisen parameters and hence are likely to be those most sensitive to anharmonic effects.
Next, we consider the phonon density of states (PDOS). In fig. 11 we show the partial phonon density of states (PDOS) computed from DFT, broken down into the contributions from each of the constituent atomic species. The advantage of computing this property is that it is intrinsically averaged across all of reciprocal space and hence offers some insight into the combined effect of many different phonon modes to the material’s thermal properties. The disadvantage is that the averaging means that information about particular phonon modes is lost.
@@
In the energy range probed by our INS experiment, the PDOS has approximately equal weight from tungsten and oxygen, but very little contribution from zirconium. This indicates that the zirconium ions are much more fixed in position, somewhat characteristic of rigid-unit-modes. Further decomposition of the PDOS down to the individual atom level (not shown) confirms that the contribution from individual tungsten ions is larger than for individual oxygen ions by about a factor two to three at low energies, but the oxygen ions are more numerous so sum up to give a comparable contribution to the PDOS.
Fig. 12 shows the PDOS with neutron weighting applied (see eq. 1), which allows direct comparison with the neutron structure factor measured in an experiment, such as that in ref. 24. The scattering lengths of all three constituent elements are similar in magnitude, as are their Debye-Waller factors, so the largest effect on the neutron cross-section comes from the division of all the other terms by the atomic mass. This gives significantly more weight to the oxygen modes than to tungsten and zirconium, with the consequence that the neutron scattering signal is dominated by oxygen motion.
@@
Figure 13: Slices through the $\mathbf{Q}$ energy plane, following the same scheme as fig. 2 of the main text. The colormaps show the neutron scattering intensity calculated using DFT, decomposed into the contributions from each individual element. Panel (a) shows the contribution from zirconium, panel (b) the contribution from tungsten, and panel (c) the contribution from oxygen. Note that the intensity scale is the same as that in fig. 3 of the main text and is the same for all three panels.